Some recent financial statements for Earl Grey Golf Corp. follow. Assets 2017 2018 Current assets Cash Accounts receivable InAsset utilization ratios: d. Total asset turnover e. Inventory turnover f. Receivables turnover times times times
Answer #1

d

Average total assets = ( beginning total assets + ending total assets)/2
Average total assets = (386581+432379)/2
Average total assets = 409480
Total asset turnover = Sales/total assets
Total asset turnover = 366996/409480
Total asset turnover = 0.9

e

Average inventory = ( beginning inventory + ending inventory)/2
Average inventory = (25392+27155)/2
Average inventory = 26273.5
Inventory turnover = COGS/inventory
Inventory turnover = 253122/26273.5
Inventory turnover = 9.63

f

Average Receivables= ( beginning Receivables+ ending Receivables)/2
Average Receivables = (12448+15235)/2
Average Receivables = 13841.5
Receivables turnover = Credit sales/receivables
Receivables turnover = 366996/13841.5
Receivables turnover = 26.51